# LinkTo-Anime: A 2D Animation Optical Flow Dataset from 3D Model Rendering

### v1

@@ 2.1 Celluloid (Cel) Animation
Celluloid animation, emerging as a form of traditional animation in the 1920s, has become a crucial foundation for the modern animation industry. [34]. In this field, animators consistently adhere to a series of systematic steps to produce cartoons, encompassing scripting, storyboarding, layout, in-between, colorization, composition, cutting,  etc. [34]. Many studies have been introduced to provide automated or supportive solutions for anime production [2, 4, 5, 10, 14, 15, 16, 31, 30, 37, 39, 42, 45]. For example, the Anime-director [14] focuses on the sequential creation of coherent story scripts and the semantics of images, contributing to automated storyboarding. On the other hand, tooncrafter [39] employs diffusion models to interpolate intermediate frames from given keyframes for in-between. Furthermore, techniques such as BasicPBC [5, 4], LVCD [10] and MangaNinja [15] have been developed to facilitate automated colorization of black-and-white line drawings, thus enhancing the effectiveness of colorization.
@@ 3.3 Data Analysis and Comparison
Table 2: Quantitative outcomes of optical flow fine-tuned on various datasets utilizing different models. ‘occ’: occluded regions; ‘s<10’: region with flow speed below 10 pixels; ‘s10-50’: region with speed ranging from 10 to 50 pixels; ‘s>50’: region with speed exceeding 50 pixels. Datasets: S: Sintel[1]; T: FlyingThing3D [17]; Cr: CreativeFlow+ [27]; R: AnimeRun [29]; LT: LinkTo-Anime.
@@
PWC-Net 119.04 20.85 119.95 126.05 3.91 13.24 119.86 4.13
PWC-Net ft. T+S 8.47 17.23 8.38 8.64 4.24 12.01 8.50 3.63
PWC-Net ft. T+Cr 0.63 7.46 0.56 0.32 5.61 13.81 0.60 4.28
PWC-Net ft. T+R 15.23 17.36 15.21 15.83 3.14 6.96 15.33 3.31
PWC-Net ft. T+LT 0.30 5.96 0.24 0.15 2.80 5.65 0.27 3.57
RAFT 9.14 19.43 9.05 9.48 3.35 11.21 9.19 3.29
RAFT ft. T+S 9.91 18.57 9.83 10.13 4.68 13.88 9.96 3.81
RAFT ft. T+Cr - - - - - -
RAFT ft. T+R 48.62 19.00 48.90 50.97 2.57 5.49 48.97 3.21
RAFT ft. T+LT 0.20 4.63 0.16 0.09 2.13 4.33 0.18 2.68
GMA 7.93 19.40 7.82 8.22 2.81 10.85 7.96 2.95
GMA ft. T+S 9.64 17.78 9.57 9.90 3.97 10.47 9.69 3.38
GMA ft. T+Cr 0.52 7.19 0.46 0.30 4.28 9.97 0.49 4.50
GMA ft. T+R 10.39 15.56 10.34 10.76 2.84 6.12 10.45 2.99
GMA ft. T+LT 0.20 4.60 0.15 0.08 2.12 4.32 0.18 2.64
GMFlow 7.09 17.41 7.00 7.33 2.80 12.09 7.12 3.20
GMFlow ft. T+S 7.93 16.13 7.86 8.18 3.24 16.45 7.96 3.49
GMFlow ft. T+Cr 0.59 6.88 0.53 0.27 5.71 14.99 0.56 4.75
GMFlow ft. T+R 13.36 20.04 13.29 14.02 2.18 8.76 13.43 2.61
GMFlow ft. T+LT 0.31 5.75 0.26 0.16 2.83 5.48 0.29 2.95
@@ 4.2 Experimental Results and Analysis
Model ft. T+LT can improve the performance. In Tab. 2, the results of the four methods demonstrate that fine-tuning on our dataset consistently improves the optical flow prediction performance for all EPE metrics. For the overall EPE, GMA ft. T + LT outperformed the second best (PWC-Net ft. T + Cr) by 0.32. RAFT ft. T + LT outperforming its original model by 8.94. PWC-Net ft. T + LT improved by 0.33 compared to its second-best (PWC-Net ft. T + Cr), and GMFlow ft. T + LT surpassed its corresponding second-best (GMFlow ft. T + Cr) by 0.28 points. The visualization results in Fig. 7 also demonstrate the superiority of our dataset, as our predictions exhibit the cleanest backgrounds and the most accurate foregrounds.
@@
However, when focusing on areas of higher motion (that is, $s\in[10,50]$ and $s>50$), the gap between model ft. T + Cr and model ft. T + LT becomes more apparent. For example, on PWC-Net, in these areas for the model fine-tuned on T + Cr is greater than when fine-tuned on Our dataset by 2.81 and 8.16, respectively. Overall, T + Cr fine-tuning tends to degrade performance in regions of high motion. Specifically, for the $s\in[10,50]$ region, the fine-tuned PWC-Net and GMA on CreativeFlow ranked last among all variants. Similarly, for the $s>50$ region, PWC-Net ft. T + Cr also yielded the worst results. This is likely because the overall motion magnitude in CreativeFlow+ [27] is relatively small. In addition, compared to our dataset, the use of simpler 3D models in CreativeFlow+ [27] can also contribute to this outcome. In contrast, although our dataset contains large ratio of blank background region, it features both large motion areas and more detailed character models. As a result, in regions with large motion($s\in[10,50]$ and $s<50$), models ft. T + LT consistently achieve the best performance in all settings, showing a clear improvement.
GMA achieves the best result. GMA fine-tuned on our dataset achieved overall the best performance, ranking first in most evaluation metrics. Its overall EPE surpassed that of the third-best method, PWC-Net, by 0.10. In the regions of $s<10$, GMA also outperforms RAFT by 0.03. This gap is relatively small, because the white background occupies comparatively large parts of the region. Both models were fine-tuned on LT, making the prediction of flows in low-motion regions (that is, where $s<10$) more accurate. Consequently, their performance in this region is comparable. In the higher-motion regions ($s\in[10,50]$ and $s>50$), GMA also showed an advantage. In the $s\in[10,50]$ range, it exceeded PWC-Net by 0.01 points and in the $s>50$ range, it exceeded PWC-Net of 0.01 points. Specifically, for the original pretrained PWC-Net, EPE reaches as high as 119.04. Upon closer inspection, we find that this unusually large error is primarily due to significant mispredictions in the small motion area ($126.05$ in “$s<10$”). This particular model weight originates from the link blow  [^6] http://content.sniklaus.com/github/pytorch-pwc/network-default.pytorch.
Analysis of Cel dataset results. We evaluate the performance of models fine-tuned under different settings on our constructed anime-industry dataset, Cel. Since obtaining full ground truth annotations for the cel dataset is challenging, we manually annotated a few key points as coarse ground truth and primarily relied on visual comparisons to assess the performance of different fine-tuning strategies. As shown in Fig. 7 (b), our dataset exhibits clear advantages in several aspects. First, compared to the pretrained network and models fine-tuned on Sintel and AnimeRun, the model fine-tuned on our dataset demonstrates significantly better performance in background regions. Second, compared to models fine-tuned on other datasets, our approach shows notable improvements in the prediction of fine details. For example, in the second column (where the character lowers their head toward the lower-left corner), our method provides a more accurate prediction of the disappearing region around the character’s chest.
@@
To investigate how our dataset can benefit cel anime-industry applications, we selected two animation production methods that utilize optical flow as a submodule. The first is BasicPBC [5], which is based on motion prediction and performs colorization of line drawings using reference frames. The second is AnimeInterp [28], which leverages bidirectional optical flow to interpolate intermediate frames from start and end frames. These two methods correspond to the Colorization and In-between stages commonly found in cel anime production pipelines. The quantitative results of both applications consistently demonstrate that our dataset improves the performance of optical flow-based Cel Animation production approaches.
@@
Anime In-between. We further investigate the impact of our dataset on optical flow-based anime in-between. To this end, we employ RAFT [35] models fine-tuned on different datasets to replace the optical flow estimation module in AnimeInterp [28], and evaluate their performance on an industry anime dataset, Cel. Specifically, we first convert the Cel dataset into a triplet format dataset like ATD-12K [28], and then generate intermediate frames from the given start and end frames using each RAFT variant. Following AnimeRun [29], we measure the interpolation quality using average PSNR and SSIM between the generated frames and the ground truth, as shown in Tab. 3(b). Compared to the baseline model using the original RAFT (PSNR: 19.16, SSIM: 0.882), only the model fine-tuned on LinkTo-Anime consistently improves both metrics. In contrast, other fine-tuned variants exhibit performance degradation to varying degrees.

### v2

@@ 2.1 Celluloid (Cel) Animation
Celluloid animation, emerging as a form of traditional animation in the 1920s, has become a crucial foundation for the modern animation industry. [34]. In this field, animators consistently adhere to a series of systematic steps to produce cartoons, encompassing scripting, storyboarding, layout, in-between, colorization, composition, cutting,  etc[34]. Many studies have been introduced to provide automated or supportive solutions for anime production [2, 4, 5, 10, 14, 15, 16, 31, 30, 37, 39, 42, 45]. For example, the Anime-director [14] focuses on the sequential creation of coherent story scripts and the semantics of images, contributing to automated storyboarding. On the other hand, tooncrafter [39] employs diffusion models to interpolate intermediate frames from given keyframes for in-between. Furthermore, techniques such as BasicPBC [5, 4], LVCD [10] and MangaNinja [15] have been developed to facilitate automated colorization of black-and-white line drawings, thus enhancing the effectiveness of colorization.
@@ 3.3 Data Analysis and Comparison
Table 2: Quantitative outcomes of optical flow fine-tuned on various datasets with different models. ‘occ’: occluded regions; ‘s<10’: region with speed below 10 pixels; ‘s10-50’: region with speed ranging from 10 to 50 pixels; ‘s>50’: region with speed exceeding 50 pixels. Datasets: S: Sintel[1]; T: FlyingThing3D [17]; Cr: CreativeFlow+ [27]; R: AnimeRun [29]; LT: LinkTo-Anime.
@@
PWC 119.04 20.85 111.81 116.49 5.14 12.95 111.92 4.55
PWC ft. T+S 8.47 18.57 8.38 8.64 4.24 12.01 8.50 3.63
PWC ft. T+Cr 0.52 6.82 0.48 0.30 4.28 9.98 0.50 4.50
PWC ft. T+R 15.23 18.74 15.21 15.83 3.14 6.96 15.33 3.31
PWC ft. T+LT 0.30 5.38 0.26 0.15 2.80 5.65 0.27 3.57
RAFT 8.48 18.49 8.40 8.61 4.88 15.68 8.52 3.82
RAFT ft. T+S 9.91 20.05 9.83 10.13 4.68 13.88 9.96 3.81
RAFT ft. T+Cr - - - - - - - -
RAFT ft. T+R 48.62 20.67 48.84 50.97 2.57 5.49 48.97 3.21
RAFT ft. T+LT 0.20 3.81 0.17 0.09 2.13 4.33 0.18 2.68
GMA 7.89 19.03 7.81 8.05 4.11 11.19 7.93 3.44
GMA ft. T+S 9.64 19.18 9.57 9.90 3.97 10.47 9.69 3.38
GMA ft. T+Cr 1.89 8.01 1.84 1.81 2.87 7.28 1.87 3.76
GMA ft. T+R 10.39 16.63 10.34 10.76 2.84 6.12 10.45 2.99
GMA ft. T+LT 0.20 3.78 0.17 0.08 2.12 4.32 0.18 2.64
GMFlow 6.47 17.60 6.38 6.51 4.45 14.70 6.49 3.73
GMFlow ft. T+S 7.24 15.48 7.18 7.27 5.20 19.57 7.27 4.11
GMFlow ft. T+Cr 0.59 6.42 0.55 0.27 5.71 14.99 0.56 4.75
GMFlow ft. T+R 12.35 20.30 12.29 12.79 3.29 8.34 12.43 3.16
GMFlow ft. T+LT 0.31 5.14 0.27 0.16 2.83 5.48 0.29 2.95
@@ 4.2 Experimental Results and Analysis
Model ft. T+LT can improve the performance. In Tab. 2, the results of the four methods demonstrate that fine-tuning on our dataset consistently improves the optical flow prediction performance for all EPE metrics. For the overall EPE, GMA ft. T + LT outperformed the second best (PWC-Net ft. T + Cr) by 0.10. RAFT ft. T + LT outperforming its original model by 8.28. PWC-Net ft. T + LT improved by 0.22 compared to its second-best (PWC-Net ft. T + Cr), and GMFlow ft. T + LT surpassed its corresponding second-best (GMFlow ft. T + Cr) by 0.28 points. The visualization results in Fig. 7 also demonstrate the superiority of our dataset, as our predictions exhibit the cleanest backgrounds and the most accurate foregrounds.
@@
However, when focusing on areas of higher motion (that is, $s\in[10,50]$ and $s>50$), the gap between model ft. T + Cr and model ft. T + LT becomes more apparent. For example, on PWC-Net, in these areas for the model fine-tuned on Our dataset is greater than model fine-tuned on T + Cr by 1.48 and 4.33, respectively. Comparing to our dataset, the use of simpler 3D models in CreativeFlow+ [27] may contribute to this outcome.
In some situation, T + R fine-tuning tends to degrade performance in regions of background area. Specifically, the fine-tuned RAFT, GMA, and GMFlow model on AnimeRun ranked last among all variants. his is likely because the background area of AnimeRun is not pure white, which is not consist to the characteristic of cel animation production images.
Analysis of Cel dataset results. We evaluate the performance of models fine-tuned under different settings on our constructed anime-industry dataset, Cel. Since obtaining full ground truth annotations for the cel dataset is challenging, we manually annotated a few key points as coarse ground truth and primarily relied on visual comparisons to assess the performance of different fine-tuning strategies. Fig. 7 (b) illustrates the distinct benefits of our dataset. Firstly, compared to models pre-trained and fine-tuned on Sintel and AnimeRun, our dataset significantly enhances model performance in background areas and fine detailed areas compared to other datasets. For instance, in the second column, where the character’s head moves downward to the lower-left, our model delivers a more precise prediction of the vanishing region near the character’s chest.
@@
(a) Visualization of line drawing colorization. The red highlighted areas indicated the incorrect colorization results.
(b) Visualization of frame interpolation on Cel dataset. Given frames 1 & 3, the intermediate frame is interpolated using AnimeInterp [28] with RAFT finetuned(ft.) on different datasets.
Figure 8: Visualization of two anime applications: (a) Anime Interpolation (b) Anime Interpolation
To examine the benefits of our dataset for cel anime-industry applications, we chose two animation production methods incorporating optical flow as a component. The first is BasicPBC [5], which employs motion prediction to colorize line drawings using reference frames. The second is AnimeInterp [28], using bidirectional optical flow for interpolating intermediate frames. Two methods align with the Colorization and In-between stages in cel anime production pipelines, respectively.
@@
Anime In-between. We study the effects of our dataset on the use of anime between. Using RAFT [35] models refined with various datasets, we substituted the optical flow module in AnimeInterp [28] and tested its efficacy in the Cel dataset. Initially, we converted the Cel data set into a triplet format, similar to ATD-12K [28], and then produced intermediate frames from 1st and 3th frames using each RAFT variant. As per AnimeRun [29], the interpolation quality is assessed via average PSNR and SSIM compared to the ground truth, as indicated in Tab. 3(b). The model fine-tuned with LT + T demonstrates an improvement in both the PSNR and SSIM metrics compared to the baseline RAFT model, which has a PSNR of 19.16 and an SSIM of 0.882. This contrasts to the performance of other fine-tuned models, which exhibit varying degrees of reduced efficacy across these metrics.